(2x+y) = 10x-4y. I need to find x and y, and the answer for both must equal 90. Of course, x and y is the same number-I need to find one number for x and one for y to make them both equal 90 degrees. Thank you!
Answer by richard1234(7193)

Wait, what you typed makes little sense. What expression equals 90 degrees? Do 2x+y and 10x-4y both equal 90 degrees? If x and y are the same number, then replace y with x in your equation to get 2x+x = 10x-4x, or 3x = 6x, which only holds when x = 0, but 2x+y and 10x-4y are not 90. If you want 2x+y = 10x-4y = 90, then we have 5y = 8x --> y = 8x/5, and you can substitute y with 8x/5 to solve for x and find y (of course, x cannot equal y in this case).